OpenStudy (anonymous):

is anyone willing to help me with a math practice ? going to take some time because there is 25 questions

OpenStudy (anonymous):

@heybhai can you help?

OpenStudy (anonymous):

@tayloranne50 if the work is not urgent you can mail me at rogerspoj@gmail.com.I will send you all with detailed solution tonight

OpenStudy (anonymous):

it has to be turned in within an hour before my teacher leaves. i need help now.

OpenStudy (anonymous):

@heybhai

OpenStudy (aripotta):

i'll try to help you, but i want you to post what you think the answers are first

OpenStudy (anonymous):

okay. the ordered pairs (0,1), (1,4), (2,16), (3,64) and (4,256) represent a function. what is a rule that represents this function? a. y = x^4 b. y = 4x c. y = x + 4 d. y = 4^x im gonna say c.

OpenStudy (anonymous):

okay @tayloranne50 send all your queries now

OpenStudy (aripotta):

c is not the answer but if you take the first point, (0,1), and put it into c, it's not true. 1 = 0 + 4? no

OpenStudy (aripotta):

because* not but

OpenStudy (anonymous):

b?

OpenStudy (aripotta):

1 = 4(0)? no

OpenStudy (anonymous):

d?

OpenStudy (aripotta):

yes, it's d lol. i hope you're not just giving out random guesses

OpenStudy (anonymous):

Suppose tuna is on sale for $0.90 a can and you have a coupon for $0.40 off your total purchase. Write a function rule for the cost of N cans of tuna A. C(n)=0.90n-0.40 B. C(n)=0.40n-0.90 C. C(n)=(0.90+n)-0.40 D. C(n)=(0.40+n)-0.90 for this one i think its b but im not positive.

OpenStudy (aripotta):

b is saying that a can is 0.40 and you have a coupon for 0.90 off...so b is backwards

OpenStudy (anonymous):

so then it has to be a?

OpenStudy (aripotta):

yes, it's a

OpenStudy (anonymous):

determine which of the mapping diagrams represents a relation that is not a function. for this one i think b but i also got confused.

OpenStudy (anonymous):

@AriPotta do you know?

OpenStudy (aripotta):

the first one doesn't pass the vertical line test, so a is not a function

OpenStudy (anonymous):

tell whether the sequence is arithmetic. if it is, what is the common difference? -12, -7, -2, 3 . . . a. yes; 5 b. yes; 8 c. yes; 3 d. no i think its d.

OpenStudy (aripotta):

it is arithmetic and the difference is 5. -12 + 5 = -7, -7 + 5 = -2, and so on

OpenStudy (anonymous):

so a?

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

id say a.

OpenStudy (aripotta):

why do you think it's a

OpenStudy (anonymous):

actually i would say d because when i multiply it doesnt come out right when i use the numbers in the other answers.

OpenStudy (aripotta):

exactly, i got that too. i'd say d as well

OpenStudy (anonymous):

what is the graph of the function rule. i have no clue about this one.

OpenStudy (aripotta):

the answer is graph d. it ends in + 2, so the vertex has to be at (0,2) and it's positive, so it's going upward, not downward

OpenStudy (anonymous):

im not sure about this one either.

OpenStudy (anonymous):

@tayloranne50 do note that mode function || creates graph in V shaped and as the value inside is +ve its upward as @AriPotta told

OpenStudy (anonymous):

For verification purpose only @tayloranne50 use graph.tk

OpenStudy (aripotta):

so for this one, we wanna convert the equation into slope-intercept form. that way, it's easier to graph. slope-intercept form is y = mx + b

OpenStudy (anonymous):

by looking at them i think its b.

OpenStudy (aripotta):

once you convert it into slope-intercept form, you get y = 3x + 4. so the slope is positive. knowing that, we can rule out choices a and b based on this

OpenStudy (aripotta):

by also looking at our equation, we see that the line has a y-intercept of 4. so it must pass through the point (0,4). so which graph shows a line with a positive slope and a y-intercept of 4?

OpenStudy (anonymous):

d?

OpenStudy (aripotta):

yes :)

OpenStudy (anonymous):

find the slope of the line. i would say b

OpenStudy (aripotta):

take 2 points off the line and use the slope formula: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] i'm going to use points (-1,3) and (0,-1) \[\frac{ 3-(-1) }{ -1 - 0 }=~\frac{ 3 + 1 }{ -1 }=~\frac{ 4 }{ -1 }=~-4\]

OpenStudy (anonymous):

what is the slope of the line that passes through the pair of points? (-2,1), (2,17) a. 4/17 b. -15 c. -4 d. 4

OpenStudy (anonymous):

im not sure about this one

OpenStudy (aripotta):

use the slope formula m8 \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

c??

OpenStudy (aripotta):

(-2,1) (2,17) \[\frac{ 17-1 }{ 2-(-2) }=~\frac{ 16 }{ 4 }=~4\]

OpenStudy (aripotta):

what are you doing to get your answers?

OpenStudy (anonymous):

i did -17 - 1 for some reason.

OpenStudy (anonymous):

i mixed it up.

OpenStudy (aripotta):

oh ok

OpenStudy (anonymous):

a. -2 b 0 c 2 d undefined id say a.

OpenStudy (aripotta):

remember that picture i uploaded?

OpenStudy (aripotta):

looking at that, and then at the graph, what do you think the answer is

OpenStudy (anonymous):

so b?

OpenStudy (aripotta):

yes, b

OpenStudy (anonymous):

im not sure about this one.

OpenStudy (aripotta):

point-slope form is y - y1 = m(x - x1). so it gives us a point (-4,6) and the slope, 3/4. put them into the formula and tell me what you get

OpenStudy (anonymous):

d?

OpenStudy (aripotta):

no, y - 6 = 3/4(x + 4) which is c

OpenStudy (anonymous):

i think d.

OpenStudy (aripotta):

no. first take two points from the table and find the slope

OpenStudy (aripotta):

after you have the slope, take one point and put the info into point-slope form. then convert it into slope-intercept form

OpenStudy (anonymous):

im not sure. i only have 15 minutes left to do this and still have like 10 problems.

OpenStudy (anonymous):

idk how to do that,

OpenStudy (aripotta):

use points (2,8) and (4,15). use the slope formula.\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]tell me what you get

OpenStudy (anonymous):

2,7?

OpenStudy (aripotta):

\[\frac{ 15-8 }{ 4-2 }=~\frac{ 7 }{ 2 }\]

OpenStudy (aripotta):

so that immediately eliminates choices c and d

OpenStudy (anonymous):

b?

OpenStudy (aripotta):

now i'm going to use (2,8) and our slope 7/2 in the point-slope form, which is y - y1 = m(x - x1) y - 8 = 7/2(x - 2) now convert it into slope-intercept form, which is y = mx + b y - 8 = 7/2x - 7 y = 7/2x + 1, which is choice a.

OpenStudy (anonymous):

write y = -1/2 x + 1 in standard form using integers. a. x + 2y = -2 b. x - 2y = -2 c. 2x + y = -2 d. x - 2y = 2 C.

OpenStudy (aripotta):

so standard form is ax + by = c where a is not negative or a fraction y = -1/2x + 1 add 1/2x to both sides 1/2x + y = 1 since a can't be a fraction, multiply both sides by 2 x + 2y = 2

OpenStudy (aripotta):

that's strange...it's not one of the choices. but i know that's the answer

OpenStudy (anonymous):

y = -1/2x + 1is what the problem is. not sure why its not one of the choices.

OpenStudy (aripotta):

i have to go walk my dog. i think you might want to get help from someone else while i'm gone. i'll probably be like 5 - 10 min